The relative fitness (w) of three genotypes are wAA=1, wAa=1, and waa=0.5, indicating that the waa genotype has fitness that is only 50% as high as the fitness of the wAA and wAa genotypes. A population starts with p=0.7 and q= 0.3, where p and q are the frequencies of the A and a alleles, respectively. What will be the value of p in the next generation, following one round of selection? Important information regarding your calculations: please provide an answer that is rounded to two decimal places, e.g., 0.69. To make calculations simpler, you should also round partial results to two decimal places (e.g., 0.675 becomes 0.68 and 0.343 becomes 0.34), as well as rounding your final answer. Credit will not be given for answers that are not in the requested format.

What is the predicted equilibrium frequency of a in the population if the mutation rate from A to a is 3x10-5? Please enter your answer in decimal format with one significant digit, for example: 0.0001 or 0.003. Credit will not be given for answers that are not in the requested format.
Answer #1

AA

Aa

aa

Initial genotype Frequency

P2= .7 X.7 = .49

2pq= 2X.7X.3 = .42

Q2= .3X.3 =.09

Fitness

1

1

.5

Genotype frequency after 1 selection

.49

.42

.045

Relative genotype frequency

.49/ (.49+.42+.045)

= .513

.42/ .955= .439

.047

The value of p after one generation= 0.51+ .44/2 = .51+.22 = 0.73

for equilibrium, rate of change due to muation will be same for both A & a.

x- 3x10-5 = (1-x) + 3x10-5

or, 2x= 1 + 6x10-5

x= 0.50003